Solve the quadratic equation by completing the square. (Enter your answers as a comma-separated list.)

x2 + 5x − 2 = 0

Found 2 solutions by lwsshak3, ewatrrr:
Answer by lwsshak3(11628)

Solve the quadratic equation by completing the square. (Enter your answers as a comma-separated list.)
x2 + 5x − 2 = 0
(x^2+5x)-2=0
To make the first expression a perfect square, take 1/2 of the coefficient(5) of the x-term and square it=(5/2)^2=25/4, then add it to the expression.
(x^2+5x+25/4)=2
since you added 25/4 to the left side, you must add 25/4 to the right side.
(x^2+5x+25/4)=2+25/4
(x+5/2)^2=33/4
take sqrt of both sides
x+5/2=±√(33/4)≈2.87
x≈-5/2±2.87
x≈-2.5±2.87
x≈-5.37
or
x≈.37
ans: x=-5.37, 0.37

Answer by ewatrrr(24785)